OpenStudy (anonymous):

which of the following equations describes the line shown below? Check all that apply. (-6,2) (-1,-4)

OpenStudy (anonymous):

Any line that travels through the points (-6,2) and (-1,-4).

OpenStudy (anonymous):

What are the equations?

OpenStudy (anonymous):

a-y-2=-5/6(x+6) b-y+4=-5/6(x+1) c-y-6=-6/5(x+2) d-y-4=-5/6(x-1) e-y-2=-6/5(x+6) f-y+4=-6/5(x+1)

OpenStudy (anonymous):

there they are

OpenStudy (anonymous):

type them in separate lines so I can make them out from one another.

OpenStudy (anonymous):

A .y-2=-5/6(x+6) B .y+4=-5/6(x+1) C .y-6=-6/5(x+2) D .y-4=-5/6(x-1) E .y-2=-6/5(x+6) F .y+4=-6/5(x+1)

OpenStudy (anonymous):

\[y - 2 = -\frac{ 5 }{ 6 }(x + 6)\] That's the format for the equations?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

im from britain and im new to this type of math

OpenStudy (anonymous):

Isolate y by adding 2 to both sides. y = -5/6(x + 6) + 2 Now you have it in slope intercept form.

OpenStudy (anonymous):

is that the only answer because it said check all that apply

OpenStudy (anonymous):

No, you still have to plug it all in.

OpenStudy (anonymous):

Okay, let me show you how.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

(-6, 2) A) y-2=-5/6(x+6) y = -5/6(x + 6) + 2 f(-6) = -5/6(-6 + 6) + 2 f(-6) = -5/6 + 2 f(-6) = 7/6 When x = -6 y = 7/6. (-6, 7/6) = (-6,2) False

OpenStudy (anonymous):

ok im starting to get it

OpenStudy (anonymous):

so a is out right

OpenStudy (anonymous):

Yeah, it's outs

OpenStudy (anonymous):

is b out

OpenStudy (anonymous):

One second, brb

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm back

OpenStudy (anonymous):

@eric1212

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

So, next one. y+4=-5/6(x+1) y = -5/6(x + 1) - 4 x = -6 y = -5/6(-6 + 1) - 4 y = -5/6 * -5 - 4 y = -5/6 * -9 y = 15/2 (-6,2) = (-6,15/2) False

OpenStudy (anonymous):

Do you see how to do it now?

OpenStudy (anonymous):

not really